Emily is 5 years older than Herman. In six years, she will be as twice as old as he is now. How old are they?

Answer by VFBundy(438)

Herman = x
Emily = x + 5

(x + 5) + 6 = 2x

x + 11 = 2x

-x = -11

x = 11

Herman = x + 11
Emily = x + 5 = 16